# HTMformer: Hybrid Time and Multivariate Transformer for Time Series Forecasting

### v2

@@ D.1 Dataset Descriptions
To ensure the consistency of our experimental approach and make sure there are no data leakage issues, we split the ETT and PEMS datasets into the training, validation, and test sets, with a 6:2:2 ratio and other datasets with a 7:1:2 ratio, following the protocols recommended in recent literature (Wu et al., 2021; Zeng et al., 2023). In all forecasting experiments, the lookback window was fixed at 96 time steps, with prediction horizons of 96, 192, 336, 720 for the long-term experiments and 12, 24, 48, 96 for the short-term experiments. The details of all datasets are provided in Table 6.
@@
  • $\ast$ The forecastability is calculated by one minus the entropy of Fourier decomposition of time series (Goerg, 2013). A larger value indicates better predictability.
@@ D.2 Baseline Details
We employ five state-of-the-art Transformer-based models, such as iTransformer (Liu et al., 2024b), PatchTST (Nie et al., 2023), FEDformer (Zhou et al., 2022), WPMixer (Murad et al., 2025), and MultiPatchFormer (Naghashi et al., 2025), as well as three well-acknowledged non-Transformer-based models, Dlinear (Zeng et al., 2023), TimeMixer (Wang et al., 2024a), and SegRNN (Lin et al., 2023). We utilized the official repositories of the Time-Series-Library (Wang et al., 2024b) directly for reproduction, and adopted the same comparison settings. The Time-Series-Library consolidates and integrates implementations of prior time-series forecasting models, making it convenient for us to conduct experiments. Among these models, MultiPatchFormer is a highly advanced architecture and yields the best overall forecasting accuracy. Both iTransformer and our model utilize input inversion to enable the attention mechanism to effectively capture multivariate correlations. WPMixer efficiently captures both the frequency-domain and time-domain information in time series data, enabling deep temporal feature extraction. FEDformer achieves high forecasting accuracy and computational efficiency through an innovative architecture that integrates seasonal-trend decomposition with frequency-domain enhancement. Meanwhile, DLinear is a state-of-the-art linear model with a non-Transformer architecture. SegRNN has demonstrated outstanding performance on multiple benchmark time series forecasting datasets by employing segment-wise iterative processing and parallelized multi-step prediction. TimeMixer is a linear time series forecasting model that leverages a multi-scale decomposable mixing mechanism, demonstrating robust performance on high-dimensional datasets. Note that SegRNN is incompatible with short-term forecasting, so no comparison is conducted.
@@ D.3 Experiment Details
All experiments were implemented in PyTorch (Paszke et al., 2019) and conducted on a dedicated NVIDIA Quadro RTX 6000 GPU. We trained the model for 10 epochs using the Adam optimizer (Kingma & Ba, 2015), with an early stopping mechanism configured at a patience value of 3. The key parameter configurations are listed in Table 7 (WPMixer is configured with d_model = 16 and d_ff = 32). We verified that, for these baselines, all other hyperparameters were taken from their official repositories within the Time-Series-Library (Wang et al., 2024b), which consolidates implementations of prior time-series forecasting models. This ensures consistency with the established fair comparison protocol. Unless otherwise stated, all models adopt the aforementioned configurations across all experiments. The only modification was to the output sequence lengths.
@@ E.1 Performance Promotion with HTME
To assess the effectiveness and scalability of the HTME for enhancing the Transformer’s capacity to learn temporal sequence representations, we integrated HTME into the Transformer and its variants: Transformer (Vaswani et al., 2017), Reformer (Kitaev et al., 2020), Informer (Zhou et al., 2021), Flowformer (Ma et al., 2023), and Flashformer (Dao et al., 2022). The length of historical data sequences was set to 96 time steps, with the forecast horizon varying among 96, 192, 336, and 720 time steps. The full predictive outcomes are detailed in Table 8. On three benchmark datasets (ECL, Traffic, and Weather), the HTME versions yield significant improvements over all four models across all proposed prediction horizons, thereby confirming their effectiveness and scalability. We further integrated the multivariate feature extraction module from the HTME extractor into iTransformer to assess the scalability capability of this module, with the detailed results presented in the ablation study. This component was seamlessly incorporated into the iTransformer’s embedding layer, leading to a substantial improvement in its overall performance.
@@
Models Transformer Reformer Informer Flowformer Flashformer
Dataset Type Step MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity Original 96 0.264 0.363 0.304 0.391 0.330 0.415 0.272 0.371 0.263 0.361
192 0.275 0.375 0.350 0.426 0.356 0.438 0.269 0.371 0.270 0.369
336 0.262 0.362 0.354 0.429 0.376 0.454 0.264 0.364 0.263 0.360
720 0.286 0.373 0.361 0.422 0.454 0.464 0.284 0.374 0.279 0.368
Avg 0.271 0.363 0.342 0.417 0.379 0.442 0.272 0.370 0.268 0.364
+HTME 96 0.157 0.247 0.169 0.266 0.170 0.266 0.170 0.266 0.163 0.262
192 0.171 0.260 0.190 0.283 0.187 0.279 0.188 0.280 0.183 0.277
336 0.192 0.277 0.200 0.293 0.207 0.296 0.202 0.294 0.199 0.292
720 0.220 0.304 0.248 0.329 0.247 0.329 0.235 0.319 0.229 0.316
Avg 0.185 0.272 0.201 0.292 0.202 0.292 0.198 0.289 0.194 0.287
Weather Original 96 0.372 0.417 0.311 0.355 0.420 0.452 0.283 0.364 0.326 0.395
192 0.632 0.580 0.453 0.464 0.423 0.455 0.548 0.541 0.587 0.544
336 0.699 0.606 0.476 0.474 0.588 0.539 0.614 0.548 0.714 0.614
720 0.903 0.688 0.579 0.541 1.146 0.784 1.079 0.783 0.906 0.684
Avg 0.651 0.572 0.454 0.458 0.644 0.557 0.631 0.559 0.633 0.559
+HTME 96 0.163 0.209 0.164 0.209 0.170 0.213 0.167 0.212 0.168 0.215
192 0.222 0.261 0.214 0.255 0.220 0.260 0.216 0.255 0.218 0.258
336 0.272 0.290 0.270 0.288 0.274 0.298 0.279 0.302 0.275 0.298
720 0.359 0.352 0.351 0.348 0.354 0.351 0.354 0.350 0.352 0.349
Avg 0.254 0.277 0.249 0.275 0.254 0.280 0.254 0.279 0.253 0.380
Traffic Original 96 0.658 0.364 0.715 0.399 0.731 0.413 0.646 0.359 0.637 0.353
192 0.655 0.359 0.705 0.390 0.761 0.431 0.646 0.355 0.659 0.364
336 0.657 0.361 0.703 0.386 0.840 0.474 0.657 0.358 0.653 0.357
720 0.697 0.380 0.703 0.383 0.969 0.545 0.671 0.364 0.697 0.375
Avg 0.666 0.366 0.706 0.389 0.825 0.465 0.655 0.359 0.661 0.362
+HTME 96 0.439 0.300 0.445 0.302 0.466 0.314 0.451 0.309 0.439 0.300
192 0.450 0.302 0.463 0.308 0.468 0.312 0.466 0.310 0.460 0.310
336 0.462 0.310 0.478 0.313 0.488 0.322 0.484 0.320 0.480 0.321
720 0.517 0.339 0.514 0.334 0.530 0.347 0.524 0.345 0.516 0.342
Avg 0.467 0.312 0.475 0.314 0.488 0.323 0.481 0.321 0.474 0.319
@@ Appendix F Additional Ablation Studies
To investigate the functional rationale of HTMformer components, we conduct detailed ablation experiments involving both component replacement and component removal. The detailed results are reported in Table 13 of the main text. As all experiments were conducted under identical configurations, the HTMformer variants are directly comparable to other models.
@@
By comparing iTransformer and HTMformerV1, blue numbers are used to denote cases where HTMformerV1 outperforms iTransformer. Observations indicate that HTMformerV1 demonstrates improvements in 63/80 cases. By comparing HTMformerV1 and HTMformerV2, numbers prefixed with an underline indicate that HTMformerV1 outperforms HTMformerV2. Notably, HTMformerV1 performs better in 57/80 cases. Significantly, on Solar and PEMS datasets lacking timestamps, HTMformerV2 often exhibits superior performance, verifying our first hypothesis: 1) The majority of time series features are stored in the temporal dimension, and the depth of temporal feature mining substantially influences prediction accuracy.
@@
Using iTransformer and HTMformerV1 as baselines, we compare them against iTransformerV3 and HTMformer versions. Red highlighted numbers indicate cases where the multivariate-integrated versions outperform their original counterparts. Experimental results show that iTransformerV3 achieves superior performance in 61/80 cases compared to iTransformer, while HTMformer outperforms its baseline in 68/80 cases. The gains are particularly evident on the PEMS datasets, which are derived from public transportation network data characterized by strong correlations among sensor nodes. This finding corroborates our second hypothesis: 2) The spatial dimension (multivariate dimension) also contains abundant features, and comprehensive multivariate feature mining can further reduce errors.
@@
Among all models and their variants, HTMformer excels in 68/80 cases. This demonstrates the effectiveness of the combination. In most cases, temporal features are the primary characteristics of time series data, but the correlations between multiple variables cannot be overlooked. It is of great significance for improving the accuracy of Transformer-based predictors. Different datasets vary in the depth of temporal and multivariate features. Therefore, we use a learnable parameter to control the proportion of the two modules, allowing the fusion of multivariate features to still reduce errors to an extent. In conclusion, the strategy of fusing temporal features and multivariate features is superior to using either of them alone. This strategy can enhance the accuracy and scalability of Transformer-based predictors.
@@
iTransV3
MPFormer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ECL 96 0.157 0.247 0.153 0.246 0.255 0.34 0.163 0.252 0.165 0.262 0.159 0.249
192 0.171 0.260 0.169 0.260 0.264 0.349 0.175 0.263 0.184 0.278 0.171 0.260
336 0.192 0.277 0.187 0.277 0.273 0.357 0.282 0.299 0.199 0.294 0.188 0.277
720 0.220 0.304 0.227 0.311 0.31 0.384 0.357 0.348 0.231 0.318 0.228 0.309
AVG 0.185 0.272 0.184 0.273 0.275 0.357 0.244 0.29 0.194 0.288 0.186 0.273
Weather 96 0.164 0.209 0.169 0.212 0.175 0.229 0.176 0.216 0.167 0.167 0.168 0.209
192 0.222 0.261 0.218 0.256 0.223 0.264 0.223 0.255 0.217 0.256 0.213 0.250
336 0.272 0.290 0.275 0.298 0.287 0.31 0.28 0.298 0.272 0.297 0.273 0.293
720 0.359 0.352 0.353 0.349 0.364 0.358 0.357 0.349 0.353 0.349 0.354 0.347
AVG 0.254 0.277 0.253 0.278 0.262 0.29 0.259 0.279 0.252 0.267 0.252 0.274
Traffic 96 0.439 0.300 0.447 0.303 0.648 0.405 0.442 0.302 0.440 0.309 0.433 0.290
192 0.450 0.302 0.460 0.307 0.737 0.438 0.459 0.308 0.457 0.305 0.450 0.296
336 0.462 0.310 0.482 0.319 0.745 0.435 0.479 0.319 0.478 0.321 0.467 0.306
720 0.517 0.339 0.522 0.343 0.863 0.484 0.516 0.342 0.520 0.349 0.500 0.326
AVG 0.467 0.312 0.477 0.318 0.748 0.44 0.474 0.317 0.473 0.321 0.462 0.304
ETTh2 96 0.300 0.348 0.298 0.348 0.375 0.399 0.295 0.344 0.304 0.352 0.299 0.350
192 0.389 0.402 0.387 0.402 0.45 0.435 0.375 0.398 0.391 0.406 0.384 0.401
336 0.409 0.408 0.416 0.427 0.481 0.469 0.436 0.435 0.419 0.428 0.411 0.428
720 0.421 0.439 0.426 0.443 0.471 0.474 0.429 0.447 0.422 0.438 0.430 0.446
AVG 0.379 0.399 0.381 0.405 0.444 0.444 0.383 0.406 0.384 0.406 0.381 0.406
Solar 96 0.199 0.239 0.209 0.252 0.244 0.284 0.213 0.253 0.207 0.249 0.203 0.240
192 0.233 0.263 0.238 0.274 0.277 0.306 0.242 0.274 0.240 0.279 0.237 0.268
336 0.252 0.279 0.261 0.292 0.323 0.326 0.262 0.29 0.256 0.290 0.249 0.280
720 0.259 0.280 0.265 0.295 0.311 0.329 0.27 0.296 0.275 0.303 0.254 0.283
AVG 0.235 0.265 0.243 0.278 0.288 0.311 0.246 0.278 0.244 0.280 0.235 0.267
PMES03 96 0.250 0.343 0.304 0.393 0.302 0.363 0.337 0.413 0.240 0.340 0.302 0.395
192 0.297 0.378 0.353 0.424 0.373 0.422 0.374 0.436 0.300 0.384 0.349 0.428
336 0.274 0.356 0.318 0.388 0.32 0.388 0.332 0.395 0.286 0.365 0.313 0.388
720 0.336 0.401 0.387 0.438 0.409 0.44 0.397 0.44 0.352 0.414 0.383 0.441
AVG 0.289 0.369 0.340 0.410 0.351 0.403 0.36 0.421 0.294 0.375 0.336 0.413
PMES04 96 0.226 0.335 0.364 0.433 0.236 0.348 0.387 0.446 0.235 0.344 0.356 0.428
192 0.286 0.383 0.406 0.406 0.326 0.413 0.429 0.473 0.307 0.395 0.398 0.459
336 0.286 0.372 0.359 0.417 0.312 0.398 0.369 0.426 0.297 0.378 0.348 0.413
720 0.338 0.414 0.436 0.471 0.387 0.451 0.44 0.476 0.362 0.428 0.423 0.467
AVG 0.284 0.376 0.391 0.431 0.315 0.402 0.406 0.455 0.300 0.386 0.381 0.441
PMES08 96 0.365 0.409 0.469 0.461 0.401 0.416 0.477 0.47 0.338 0.388 0.406 0.437
192 0.513 0.462 0.612 0.507 0.567 0.481 0.627 0.518 0.487 0.444 0.539 0.479
336 0.530 0.435 0.606 0.471 0.567 0.452 0.599 0.472 0.516 0.425 0.538 0.444
720 0.624 0.487 0.692 0.531 0.697 0.512 0.691 0.532 0.600 0.479 0.603 0.489
AVG 0.508 0.448 0.594 0.492 0.558 0.465 0.598 0.498 0.485 0.434 0.521 0.462

### v3

@@ D.1 Dataset Descriptions
To ensure the consistency of our experimental approach and make sure there are no data leakage issues, we split the ETT and PEMS datasets into the training, validation, and test sets, with a 6:2:2 ratio and other datasets with a 7:1:2 ratio, following the protocols recommended in recent literature (wu2021autoformer; zeng2023dlinear). In all forecasting experiments, the lookback window was fixed at 96 time steps, with prediction horizons of 96, 192, 336, 720 for the long-term experiments and 12, 24, 48, 96 for the short-term experiments. The details of all datasets are provided in Table 5.
@@
  • $\ast$ The forecastability is calculated by one minus the entropy of Fourier decomposition of time series (goerg2013forecastable). A larger value indicates better predictability.
@@ D.2 Baseline Details
We employ four state-of-the-art Transformer-based models, such as iTransformer (liu2024itransformer), PatchTST (nie2023time), FEDformer (zhou2022fedformer), and MultiPatchFormer (naghashi2025multipatchformer), two well-acknowledged Linear models, Dlinear (zeng2023dlinear), Nlinear, one prominent RNN-based model SegRNN (lin2023segrnn) and WPMixer (murad2024wp). We comprehensively evaluate the HTMformer framework using four mainstream attention mechanisms Transformer (vaswani2017attention), Reformer (kitaev2020reformer), Informer (zhou2021informer), Flowformer (ma2023flowformer), and Flashformer (dao2022flashattention). We utilized the official repositories of the Time-Series-Library (wang2024tssurvey) directly for reproduction, and adopted the same comparison settings. The Time-Series-Library consolidates and integrates implementations of prior time-series forecasting models, making it convenient for us to conduct experiments. Among these models, MultiPatchFormer is a highly advanced architecture and yields the best overall forecasting accuracy. iTransformer and HTMformers share the same architecture. HTMformers replaces the embedding layer of iTransformer. Note that SegRNN is incompatible with short-term forecasting, so no comparison is conducted.
@@ D.3 Experiment Details
All experiments were implemented in PyTorch (paszke2019pytorch) and conducted on a dedicated NVIDIA Quadro RTX 6000 GPU. We trained the model for 10 epochs using the Adam optimizer (kingma2015adam), with an early stopping mechanism configured at a patience value of 3. The key parameter configurations are listed in Table 6. We verified that, for these baselines, all other hyperparameters were taken from their official repositories within the Time-Series-Library (wang2024tssurvey), which consolidates implementations of prior time-series forecasting models. This ensures consistency with the established fair comparison protocol. Unless otherwise stated, all models adopt the aforementioned configurations across all experiments. The only modification was to the output sequence lengths.
@@ E.1 Performance Promotion with HTME
To assess the effectiveness and scalability of the HTMformers for enhancing the Attention’s capacity to learn temporal sequence representations, we integrated various attention variants: Transformer (vaswani2017attention), Reformer (kitaev2020reformer), Informer (zhou2021informer), Flowformer (ma2023flowformer), and Flashformer (dao2022flashattention). We apply the HTME strategy to models with other architectures, including the Transformer-based iTransformer, the linear-based DLinear, NLinear and the RNN-based SegRNN. The length of historical data sequences was set to 96 time steps, with the forecast horizon varying among 96, 192, 336, and 720 time steps. The full predictive outcomes are detailed in Table 7. On three benchmark datasets (ECL, Traffic, and Weather), the HTME versions yield significant improvements over all four attentions across all proposed prediction horizons. Across three mainstream forecasting architectures, the HTME strategy consistently improves performance by margins ranging from 1.7% to 4.6%. These results highlight the effectiveness and robustness of the HTME strategy.
@@
Models Transformer Reformer Informer Flowformer Flashformer iTransformer DLinear Nlinear SegRNN
Dataset Type Step MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Electricity Original 96 0.264 0.363 0.304 0.391 0.330 0.415 0.272 0.371 0.263 0.361 0.163 0.252 0.210 0.301 0.199 0.277 0.168 0.264
192 0.275 0.375 0.350 0.426 0.356 0.438 0.269 0.371 0.270 0.369 0.175 0.263 0.210 0.304 0.199 0.280 0.183 0.279
336 0.262 0.362 0.354 0.429 0.376 0.454 0.264 0.364 0.263 0.360 0.282 0.299 0.223 0.319 0.214 0.295 0.192 0.289
720 0.286 0.373 0.361 0.422 0.454 0.464 0.284 0.374 0.279 0.368 0.357 0.348 0.257 0.349 0.256 0.328 0.231 0.321
Avg 0.271 0.363 0.342 0.417 0.379 0.442 0.272 0.370 0.268 0.364 0.244 0.290 0.225 0.318 0.217 0.295 0.194 0.288
+HTME 96 0.157 0.247 0.169 0.266 0.170 0.266 0.170 0.266 0.163 0.262 0.157 0.247 0.204 0.299 0.189 0.273 0.164 0.258
192 0.171 0.260 0.190 0.283 0.187 0.279 0.188 0.280 0.183 0.277 0.171 0.260 0.211 0.307 0.192 0.273 0.175 0.269
336 0.192 0.277 0.200 0.293 0.207 0.296 0.202 0.294 0.199 0.292 0.192 0.277 0.222 0.318 0.206 0.290 0.192 0.285
720 0.220 0.304 0.248 0.329 0.247 0.329 0.235 0.319 0.229 0.316 0.220 0.304 0.24 0.331 0.246 0.329 0.230 0.320
Avg 0.185 0.272 0.201 0.292 0.202 0.292 0.198 0.289 0.194 0.287 0.185 0.272 0.219 0.314 0.208 0.291 0.190 0.283
Weather Original 96 0.372 0.417 0.311 0.355 0.420 0.452 0.283 0.364 0.326 0.395 0.176 0.216 0.193 0.233 0.196 0.237 0.168 0.23
192 0.632 0.580 0.453 0.464 0.423 0.455 0.548 0.541 0.587 0.544 0.223 0.255 0.236 0.268 0.241 0.279 0.215 0.277
336 0.699 0.606 0.476 0.474 0.588 0.539 0.614 0.548 0.714 0.614 0.280 0.298 0.288 0.304 0.293 0.317 0.274 0.319
720 0.903 0.688 0.579 0.541 1.146 0.784 1.079 0.783 0.906 0.684 0.353 0.346 0.359 0.349 0.366 0.361 0.359 0.378
Avg 0.651 0.572 0.454 0.458 0.644 0.557 0.631 0.559 0.633 0.559 0.259 0.279 0.269 0.289 0.274 0.299 0.254 0.301
+HTME 96 0.163 0.209 0.164 0.209 0.170 0.213 0.167 0.212 0.168 0.215 0.163 0.209 0.177 0.223 0.168 0.230 0.161 0.226
192 0.222 0.261 0.214 0.255 0.220 0.260 0.216 0.255 0.218 0.258 0.222 0.261 0.227 0.266 0.224 0.275 0.211 0.273
336 0.272 0.290 0.270 0.288 0.274 0.298 0.279 0.302 0.275 0.298 0.272 0.290 0.281 0.303 0.295 0.313 0.269 0.315
720 0.359 0.352 0.351 0.348 0.354 0.351 0.354 0.350 0.352 0.349 0.359 0.352 0.358 0.352 0.359 0.343 0.355 0.371
Avg 0.254 0.277 0.249 0.275 0.254 0.280 0.254 0.279 0.253 0.380 0.254 0.277 0.261 0.286 0.261 0.290 0.249 0.296
Traffic Original 96 0.658 0.364 0.715 0.399 0.731 0.413 0.646 0.359 0.637 0.353 0.442 0.302 0.696 0.428 0.654 0.396 0.723 0.376
192 0.655 0.359 0.705 0.390 0.761 0.431 0.646 0.355 0.659 0.364 0.459 0.308 0.646 0.407 0.607 0.374 0.737 0.385
336 0.657 0.361 0.703 0.386 0.840 0.474 0.657 0.358 0.653 0.357 0.479 0.319 0.653 0.409 0.615 0.376 0.766 0.405
720 0.697 0.380 0.703 0.383 0.969 0.545 0.671 0.364 0.697 0.375 0.516 0.342 0.694 0.428 0.654 0.417 0.818 0.423
Avg 0.666 0.366 0.706 0.389 0.825 0.465 0.655 0.359 0.661 0.362 0.474 0.317 0.672 0.418 0.632 0.391 0.761 0.397
+HTME 96 0.439 0.300 0.445 0.302 0.466 0.314 0.451 0.309 0.439 0.300 0.439 0.300 0.626 0.393 0.624 0.394 0.696 0.366
192 0.450 0.302 0.463 0.308 0.468 0.312 0.466 0.310 0.460 0.310 0.450 0.302 0.583 0.369 0.588 0.372 0.717 0.382
336 0.462 0.310 0.478 0.313 0.488 0.322 0.484 0.320 0.480 0.321 0.462 0.310 0.591 0.373 0.597 0.376 0.754 0.401
720 0.517 0.339 0.514 0.334 0.530 0.347 0.524 0.345 0.516 0.342 0.517 0.339 0.636 0.395 0.635 0.395 0.803 0.419
Avg 0.467 0.312 0.475 0.314 0.488 0.323 0.481 0.321 0.474 0.319 0.467 0.312 0.609 0.383 0.611 0.384 0.750 0.392
@@ E.2 Long-term Forecasting Results
nie2023time
zeng2023dlinear
zhou2022fedformer
naghashi2025multipatchformer
murad2024wp
wang2023timemixer
lin2023segrnn
@@ E.3 Short-term Forecasting Results
nie2023time
zeng2023dlinear
zhou2022fedformer
naghashi2025multipatchformer
murad2024wp
wang2023timemixer
@@ Appendix F Additional Ablation Studies
To investigate the functional rationale of HTME components, We examined three different variants of iTransformer, DLinear, and SegRNN, as representatives of Transformer-based, Linear-based, and RNN-based models, respectively. 1) Productor, which integrates the multivariate feature extraction module of the HTME extractor into the embedding layer of iTransformer. 2) ProductorV1, which utilizes only the multivariate feature extraction module of the HTME extractor. We also examined another variant 3) ProductorV2, which utilizes the multivariate feature extraction module of the HTME extractor to supplement the embedding representations. We introduce two Additional variants: HTMformer, builds upon iTransformer by incorporating the complete HTME extractor; HTMformerV1, replaces the original embedding layer in iTransformer with the temporal feature extractor from HTME. The detailed results are reported in Table 12 of the main text. As all experiments were conducted under identical configurations, the HTMPredictors are directly comparable to other models.
@@
By comparing iTransformer and HTMformerV1, blue numbers are used to denote cases where HTMformerV1 outperforms iTransformer. Observations indicate that HTMformerV1 demonstrates improvements in 63/80 cases.
Effect of the temporal feature.
By comparing Predictor and PredictorV2, numbers prefixed with an underline indicate that Predictor outperforms PredictorV2. Notably, Predictor performs better in most cases. The majority of time series features are stored in the temporal dimension, and the depth of temporal feature mining substantially influences prediction accuracy. This indicates that the temporal feature extraction module of HTME is more effective.
@@
Using Predictor as baselines, we compare them against PredictorV1. Red highlighted numbers indicate cases where the multivariate-integrated versions outperform their original counterparts. Experimental results show that PredictorV1 achieves superior performance most cases. We also find that PredictorV2 outperforms Predictor in some cases. The spatial dimension (multivariate dimension) also contains abundant features, and comprehensive multivariate feature mining can further reduce errors.
@@
Among all models and their variants, PredictorV1 achieves superior performance. This demonstrates the effectiveness of the HTME strategy. In most cases, temporal features are the primary characteristics of time series data, but the correlations between multiple variables cannot be overlooked. The HTME strategy has successfully suppressed noise and extracted effective patterns from the raw sequence. In conclusion, the strategy of fusing temporal features and multivariate features can enhance the accuracy and scalability of Transformer-based predictors.
@@
iTransV1
DLinearV2
DLinear
DLinearV1
SegRNNV2
SegRNN
SegRNNV1
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ECL 96 0.157 0.247 0.153 0.246 0.255 0.34 0.163 0.252 0.165 0.262 0.499 0.533 0.210 0.301 0.204 0.299 0.390 0.454 0.168 0.264 0.164 0.258
192 0.171 0.260 0.169 0.260 0.264 0.349 0.175 0.263 0.184 0.278 0.505 0.535 0.210 0.304 0.211 0.307 0.407 0.464 0.183 0.279 0.175 0.269
336 0.192 0.277 0.187 0.277 0.273 0.357 0.282 0.299 0.199 0.294 0.504 0.531 0.223 0.319 0.222 0.318 0.415 0.470 0.192 0.289 0.192 0.285
720 0.220 0.304 0.227 0.311 0.31 0.384 0.357 0.348 0.231 0.318 0.491 0.526 0.257 0.349 0.240 0.331 0.452 0.491 0.231 0.321 0.230 0.320
AVG 0.185 0.272 0.184 0.273 0.275 0.357 0.244 0.29 0.194 0.288 0.500 0.531 0.225 0.318 0.219 0.314 0.416 0.470 0.194 0.288 0.190 0.283
Weather 96 0.164 0.209 0.169 0.212 0.175 0.229 0.176 0.216 0.167 0.167 0.18 0.272 0.193 0.233 0.177 0.223 0.165 0.231 0.168 0.230 0.161 0.226
192 0.222 0.261 0.218 0.256 0.223 0.264 0.223 0.255 0.217 0.256 0.232 0.320 0.236 0.268 0.227 0.266 0.221 0.285 0.215 0.277 0.211 0.273
336 0.272 0.290 0.275 0.298 0.287 0.31 0.28 0.298 0.272 0.297 0.294 0.365 0.288 0.304 0.281 0.303 0.281 0.330 0.274 0.319 0.269 0.315
720 0.359 0.352 0.353 0.349 0.364 0.358 0.357 0.349 0.353 0.349 0.391 0.433 0.359 0.349 0.358 0.352 0.364 0.385 0.359 0.378 0.355 0.371
AVG 0.254 0.277 0.253 0.278 0.262 0.29 0.259 0.279 0.252 0.267 0.274 0.348 0.269 0.289 0.261 0.286 0.258 0.308 0.254 0.301 0.249 0.296
Traffic 96 0.439 0.300 0.447 0.303 0.648 0.405 0.442 0.302 0.440 0.309 0.955 0.544 0.696 0.428 0.626 0.393 1.224 0.583 0.723 0.376 0.696 0.366
192 0.450 0.302 0.460 0.307 0.737 0.438 0.459 0.308 0.457 0.305 0.970 0.543 0.646 0.407 0.583 0.369 1.201 0.568 0.737 0.385 0.717 0.382
336 0.462 0.310 0.482 0.319 0.745 0.435 0.479 0.319 0.478 0.321 0.982 0.550 0.653 0.409 0.591 0.373 1.285 0.595 0.766 0.405 0.754 0.401
720 0.517 0.339 0.522 0.343 0.863 0.484 0.516 0.342 0.520 0.349 0.997 0.548 0.694 0.428 0.636 0.395 1.343 0.614 0.818 0.423 0.803 0.419
AVG 0.467 0.312 0.477 0.318 0.748 0.44 0.474 0.317 0.473 0.321 0.976 0.546 0.672 0.418 0.609 0.383 1.263 0.590 0.761 0.397 0.750 0.392
ETTh2 96 0.300 0.348 0.298 0.348 0.375 0.399 0.295 0.344 0.304 0.352 0.357 0.409 0.341 0.395 0.343 0.399 0.339 0.389 0.295 0.355 0.292 0.345
192 0.389 0.402 0.387 0.402 0.45 0.435 0.375 0.398 0.391 0.406 0.497 0.491 0.481 0.479 0.480 0.477 0.417 0.431 0.375 0.397 0.371 0.396
336 0.409 0.408 0.416 0.427 0.481 0.469 0.436 0.435 0.419 0.428 0.610 0.553 0.592 0.542 0.597 0.540 0.476 0.476 0.414 0.436 0.416 0.430
720 0.421 0.439 0.426 0.443 0.471 0.474 0.429 0.447 0.422 0.438 0.816 0.658 0.840 0.661 0.820 0.652 0.474 0.489 0.443 0.469 0.436 0.462
AVG 0.379 0.399 0.381 0.405 0.444 0.444 0.383 0.406 0.384 0.406 0.570 0.528 0.563 0.519 0.560 0.517 0.427 0.446 0.382 0.414 0.379 0.408
Solar 96 0.199 0.239 0.209 0.252 0.244 0.284 0.213 0.253 0.207 0.249 0.231 0.326 0.289 0.377 0.225 0.317 0.353 0.416 0.232 0.294 0.203 0.291
192 0.233 0.263 0.238 0.274 0.277 0.306 0.242 0.274 0.240 0.279 0.262 0.348 0.319 0.397 0.260 0.342 0.404 0.448 0.253 0.306 0.230 0.303
336 0.252 0.279 0.261 0.292 0.323 0.326 0.262 0.29 0.256 0.290 0.302 0.373 0.352 0.415 0.296 0.367 0.418 0.455 0.264 0.310 0.257 0.321
720 0.259 0.280 0.265 0.295 0.311 0.329 0.27 0.296 0.275 0.303 0.313 0.378 0.356 0.412 0.309 0.373 0.482 0.516 0.260 0.306 0.281 0.349
AVG 0.235 0.265 0.243 0.278 0.288 0.311 0.246 0.278 0.244 0.280 0.277 0.356 0.329 0.400 0.273 0.350 0.414 0.459 0.252 0.304 0.243 0.316
PMES03 96 0.250 0.343 0.304 0.393 0.302 0.363 0.337 0.413 0.240 0.340 0.315 0.418 0.458 0.517 0.313 0.416 0.389 0.446 0.381 0.438 0.287 0.374
192 0.297 0.378 0.353 0.424 0.373 0.422 0.374 0.436 0.300 0.384 0.410 0.477 0.477 0.527 0.410 0.477 0.620 0.602 0.447 0.478 0.368 0.433
336 0.274 0.356 0.318 0.388 0.32 0.388 0.332 0.395 0.286 0.365 0.337 0.422 0.396 0.456 0.344 0.420 0.729 0.641 0.380 0.432 0.349 0.418
720 0.336 0.401 0.387 0.438 0.409 0.44 0.397 0.44 0.352 0.414 0.391 0.469 0.439 0.492 0.396 0.462 0.774 0.683 0.447 0.478 0.406 0.472
AVG 0.289 0.369 0.340 0.410 0.351 0.403 0.36 0.421 0.294 0.375 0.363 0.447 0.442 0.498 0.366 0.444 0.628 0.593 0.413 0.456 0.353 0.424
PMES04 96 0.226 0.335 0.364 0.433 0.236 0.348 0.387 0.446 0.235 0.344 0.315 0.404 0.452 0.504 0.312 0.401 0.370 0.457 0.415 0.468 0.291 0.398
192 0.286 0.383 0.406 0.406 0.326 0.413 0.429 0.473 0.307 0.395 0.377 0.451 0.477 0.527 0.372 0.451 0.723 0.665 0.496 0.510 0.463 0.507
336 0.286 0.372 0.359 0.417 0.312 0.398 0.369 0.426 0.297 0.378 0.330 0.411 0.396 0.456 0.330 0.410 0.789 0.694 0.420 0.461 0.453 0.508
720 0.338 0.414 0.436 0.471 0.387 0.451 0.44 0.476 0.362 0.428 0.401 0.460 0.439 0.492 0.400 0.460 0.850 0.729 0.479 0.502 0.714 0.677
AVG 0.284 0.376 0.391 0.431 0.315 0.402 0.406 0.455 0.300 0.386 0.356 0.432 0.441 0.494 0.354 0.431 0.683 0.636 0.452 0.485 0.480 0.522
PMES08 96 0.365 0.409 0.469 0.461 0.401 0.416 0.477 0.47 0.338 0.388 0.449 0.446 0.672 0.564 0.454 0.451 0.429 0.477 0.516 0.487 0.364 0.425
192 0.513 0.462 0.612 0.507 0.567 0.481 0.627 0.518 0.487 0.444 0.599 0.536 0.726 0.580 0.597 0.535 0.811 0.674 0.646 0.528 0.568 0.523
336 0.530 0.435 0.606 0.471 0.567 0.452 0.599 0.472 0.516 0.425 0.569 0.476 0.650 0.518 0.564 0.469 0.949 0.712 0.617 0.490 0.621 0.511
720 0.624 0.487 0.692 0.531 0.697 0.512 0.691 0.532 0.600 0.479 0.633 0.531 0.713 0.562 0.632 0.521 1.034 0.743 0.707 0.541 0.874 0.652
AVG 0.508 0.448 0.594 0.492 0.558 0.465 0.598 0.498 0.485 0.434 0.563 0.497 0.690 0.556 0.562 0.494 0.806 0.652 0.621 0.511 0.607 0.528